When an unbiased coin is tossed seven times, what is the probability of obtaining at least 3 tails?


1
Expert's answer
2021-05-07T09:46:49-0400

Since the coin is unbiased, probabilities of heads and tails are equal. We denote by "X" a random variable that corresponds to the number of tails. It has a binomial distribution with parameters 7 and 0.5. The aim is to find the probability "P(X\geq3)". We have: "P(X\geq3)=1-P(X<3)=1-P(X=0)-P(X=1)-P(X=2)=1-(0.5)^7-C_{7}^1(0.5)^7-C_{7}^2(0.5)^7=1-(0.5)^7-7(0.5)^7-\frac{6\cdot7}{2}(0.5)^7=1-29(0.5)^7=1-29\cdot0.0078125\approx0.773"